if you have 405 apples what will be a 5% increase
Answer by Alan3354(69443)

= 405*1.05 = 425.25 apples
Or, 5% = 0.05
405*0.05 = 20.25 apples
405 + 20.25 = 425.25
--------------------
worms ate the missing 0.75 of an apple.